An integer is four less than twice another. If the product of the two integers is 70, one pair of integers that will satisfy the problem is _________
Answer by CubeyThePenguin(3113)

integers: x, y

x = 2y - 4
xy = 70

Substitute the first equation into the second.

(2y - 4)(y) = 70
2y^2 - 4y = 70
y^2 - 2y = 35
y^2 - 2y - 35 = 0
(y - 7)(y + 5) = 0

The integers are (x, y) = (-14, -5) and (10, 7).